Amsaca Products is a decentralized wholesaler with four autonomous divisions. The divisions are evaluated on the basis of ROI, with year-end bonuses given to divisional managers who have the highest ROI. Operating results for the company’s East Division for last year are given below:

Last Year

New Product Line

Last Year + New Product Line

Sales

$21,000,000

$9,000,000

$30,000,000

Variable expenses

$13,400,000

Contribution margin

$ 7,600,000

Fixed expenses

$ 5,920,000

Net operating income

$ 1,680,000

Operating assets

$ 5,250,000

The company had an overall ROI of 18% last year (considering all divisions). The company’s East Division has an opportunity to add a new product line that would require an investment in operating assets of $3,000,000. The cost and revenue characteristics of the new product line per year would be as follows:

Sales                            $9,000,000

Variable expenses       65% of sales

Fixed expenses           $2,520,000

Complete the table above for the new product line and then add the new product line to the results from last year.

  1. Compute the East Division’s ROI for last year as given above in terms of margin and turnover.
  1. Compute the ROI on the new product line in terms of margin and turnover.
  1. Compute the East Division’s ROI in terms of margin and turnover as it would appear if the new product line is added.
  1. If you were in Tracey Hawkins’ position, would you accept or reject the new product line? Explain
  1. Why do you suppose headquarters is eager for the East Division to add the new product line?
  1. Suppose that the company’s minimum required rate of return on operating assets is 15% and that performance is evaluated using residual income.
  1. Compute the East Division’s residual income for last year, using the table below. (I did that for you.)
  1. Compute the residual income on the new product line, using the table below.
  1. Compute the East Division’s residual income as it would appear if the new product line is added, using the table below.

Last Year

New Product Line

Last Year + New Product Line

Operating assets

$5,250,000

Net operating income

$1,680,000

Minus the minimum required return on operating assets

$787,500

Residual income

$892,500

  1. Under these circumstances, if you were in Tracey Hawkins’ position, would you accept or reject the new product line? Explain.
Answer #1

Last Year

New Product Line

Last Year + New Product Line

Sales

$21,000,000

$9,000,000

$30,000,000

Variable expenses

$13,400,000

5,850,000

19,250,000

Contribution margin

$ 7,600,000

3,150,000

10,750,000

Fixed expenses

$ 5,920,000

2,520,000

8,440,000

Net operating income

$ 1,680,000

630,000

2,310,000

Operating assets

$ 5,250,000

3,000,000

8,250,000

East Division ROI:

Margin = Net income/Sales = 1,680,000/21,000,000 = 8%

Turnover = Sales/Operating Assets = 21,000,000/5,250,000 = 4

Hence, ROI = Margin*Turnover = 8%*4 = 32%

New Product Line:

Margin = 630,000/9,000,000 = 7%

Turnover = 9,000,000/3,000,000 = 3

ROI = 7%*3 = 21%

When new product line is added

Margin = 2,310,000/30,000,000 = 7.7%

Turnover = 30,000,000/8,250,000 = 3.6363 times

ROI = 7.7%*3.6363 = 28%

REJECT, since ROI will reduce and bonuses are based on ROI

Yes, because ROI of new division is higher than overall ROI of 18%

RI

Last Year

New Product Line

Last Year + New Product Line

Operating assets

$5,250,000

3,000,000

8,250,000

Net operating income

$1,680,000

630,000

2,310,000

Minus the minimum required return on operating assets

$787,500

450,000

1,237,500

Residual income

$892,500

180,000

1,072,500

ACCEPT, since positive residual income